OpenStudy (anonymous):

Help please :o) Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = x-7/x+3 and g(x) = -3x-7/x-1

OpenStudy (anonymous):

\[f(x)=\frac{ x-7 }{x+3 } ; g(x)=\frac{ -3x-7 }{ x-1 }\]

zepdrix (zepdrix):

\[\large \color{#CC0033}{f(x)=\frac{x-7}{x+3}};\qquad \color{#3366CF}{g(x)=\frac{-3x-7}{x-1}}\] \[\large \color{#CC0033}{f(\color{#3366CF}{g(x)})}=\color{#CC0033}{\frac{\color{#3366CF}{g(x)}-7}{\color{#3366CF}{g(x)}+3}}\] This is going to be rather messy plugging this all in :) lol\[\huge =\color{#CC0033}{\frac{\color{#3366CF}{\frac{-3x-7}{x-1}}-7}{\color{#3366CF}{\frac{-3x-7}{x-1}}+3}}\]

zepdrix (zepdrix):

They want you to simplify it from there? D: Oh boy....

OpenStudy (anonymous):

I think I figured it out: f(x)= 0=x-7/x+3 x= -7/3 y=-7-7/-3+3 y=0 (0, -7/3) g(x)= 0=-3x-7/x-1 0=-7/-2x-1 0=-7/-3 (-7/3, 0) ??? Maybe haha

zepdrix (zepdrix):

Hmm <:o

OpenStudy (anonymous):

Thanks for trying to help me though, I appreciate it :o)